OpenStudy (anonymous):

The function f(x) = 4(4)^x represents the growth of a fly population every year in a remote swamp. Jackie wants to manipulate the formula to an equivalent form that calculates three times a year, not just once a year. Which function is correct for Jackie's purpose, and what is the new growth rate?

OpenStudy (anonymous):

@dtan5457

OpenStudy (anonymous):

can you help me with this step by step?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what does x represent?

OpenStudy (anonymous):

x=times during the year?

OpenStudy (anonymous):

so in this case x=3 ? right?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

x is the number of years

jimthompson5910 (jim_thompson5910):

so if x = 3, then that represents year 3

OpenStudy (anonymous):

so i would plant this like: f(3)=4(4)^3 ?

jimthompson5910 (jim_thompson5910):

how do we adjust things so it calculates 3 times a year

jimthompson5910 (jim_thompson5910):

nope

OpenStudy (anonymous):

we multiply by 3 so they would adjust?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes x would turn into 3x

OpenStudy (anonymous):

so the function would now be f(x)=4(4)^3x ?

jimthompson5910 (jim_thompson5910):

yes \[\Large f(x) = 4(4)^{3x}\]

OpenStudy (anonymous):

so i would solve this function?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what do you mean?

OpenStudy (anonymous):

oh wait nevermind lol we dont solve it . i just saw my options lol

OpenStudy (anonymous):

so would f(x)=4(4)^3x would be my final answer? amd the rate percentage would be 4% ?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

can i see a screenshot of all the options?

OpenStudy (anonymous):

yes here it is

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

ok let me think

OpenStudy (anonymous):

ok , to find the rate id think of graphing it, but when i graph its like a really straight line

jimthompson5910 (jim_thompson5910):

\[\Large e^{kx} = 4^{3x}\] \[\Large \ln(e^{kx}) = \ln(4^{3x})\] \[\Large kx*\ln(e) = 3x*\ln(4)\] \[\Large k*1 = 3*\ln(4)\] \[\Large k = 3*\ln(4)\] \[\Large k \approx 4.15888\] So the growth rate is approximately 4%

jimthompson5910 (jim_thompson5910):

so yes, it is B

OpenStudy (anonymous):

what was the formula you used to solve for the rate? is it the average rate of change formula?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

I set e^(kx) equal to 4^(3x) and solved for k

OpenStudy (anonymous):

thank you so much :)

jimthompson5910 (jim_thompson5910):

np